The product of two improper fractions is less than both the fractions.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is False.
The product of two improper fractions is more than each of the fractions that are multiplied.

Consider the two improper fractions, 5/4 and 4/2
= (5/4) × (4/2)

= 5/2

= 2.5
Then, 2.5 is multiplied to the improper fraction more than each of the fractions that are multiplied = (5/4) × 2.5
= 3.125
And (4/2) × 2.5 = 5

Thanks